# Stochastic activations

### v1

@@ 1 Introduction
The majority of the LLM parameters are in the Feed-Forward Network (FFN) layers, where they memorize the training data. FFNs are two linear layers separated by an activation function, and sometimes an additional linear layer that serves as a gating operation. The activation is a $\mathbb{R}\rightarrow\mathbb{R}$ non-linear function. In this context, the choice of activation function plays a crucial role for both the model’s expressivity and efficiency. The simplest activation is RELU (Rectified Linear Unit), that allows positive values to pass through and forces negative inputs to zero. RELU is sparsity-inducing, since, on average, half of its outputs are zero (in practice significantly more). Within a two-layer Multilayer Perceptron, this means that inference on the second layer is a matrix-sparse vector multiplication, so it can be implemented with fewer FLOPs than a matrix-dense vector multiplication. Note that effectively improving the runtime with this sparsity pattern remains challenging.
@@
In this work, we consider two ways to approach this problem. The first approach is activation fine-tuning, denoted Swi+FT: we pre-train the model with an activation that facilitates efficient large language model optimization, then we change the activation to RELU and adapt the model by fine-tuning it further. Our second approach, referred to as StochA (stochastic activations), is a novel technique that randomly selects between multiple activations, either at train or test time. Both approaches allow models to benefit from the superior optimization properties of SILU. These hybrid strategies combine the best of both worlds – maintaining high model performance while unlocking the computational efficiency of sparse activations.
@@
  • We introduce and analyze two strategies that employ different activation functions at training and inference time, namely Swi+FT and StochA. Both are complementary and make it possible to use activations at inference time that differ from those employed during pre-training.
@@
  • We show that stochastic activations, when used at inference time, provides an alternative way to generate diverse sequences compared to traditional temperature sampling or other variants.
@@
Activations are at the core of deep learning, in that they are needed to depart from defining a linear function with limited expressivity. While early neural network architectures were inspired by logistic regression, such as sigmoidal and tanh activations, many activation functions have been evaluated for the Feed Forward layers (FFNs) of transformers. Vaswani et al. (2017) used RELU (Glorot et al., 2011). However, using the RELU activation function leads to some neurons getting stuck in the negative region. As a consequence, they stop learning entirely, since the gradient is zero for negative inputs, and their weights do not get updated. In contrast, Touvron et al. (2023) used SILU as the activation function for the FFN layers of the transformer for the first Llama models. Shazeer (2020) discusses the benefits of SWIGLU, which consists of SILU with gating. There exist many other activation functions such as the Gaussian Error Linear Unit (GELU) (Hendrycks & Gimpel, 2016), Scaled Exponential Linear Unit (SELU) (Klambauer et al., 2017), Swish (Ramachandran et al., 2018) and gated SILU, among others.
@@
Lee et al. (2022) propose an activation, called Adaptive Swish (ASH), that uses stochastic sampling of the top-k percentile elements. It is a generalization of the Swish (Ramachandran et al., 2018) activation function, that uses adaptive thresholding and selects only the values in the top percentiles and sets these to zero otherwise. The threshold is not fixed but depends on the distribution of the input making it an example of a stochastic activation.
@@
LayerDrop (Fan et al., 2019) randomly drops entire layers during training, hence, it encourages the model to be robust to missing layers. At inference time, some layers can be pruned, trading off between speed and accuracy as needed. While the method does not make the model sparse in the usual sense, it induces structured sparsity in the computation graph during training. Other works also introduce structured dropout variants such as DropBlock (Ghiasi et al., 2018), Bayesian dropout (Gal & Ghahramani, 2016), or Beit (Bao et al., 2021) and masked-autoencoder (He et al., 2022) in computer vision, among others.
@@ 3.1 Fine-tuning with RELU: Swi+FT
Figure 2: Swi+FT: Training loss. Most of the training is carried out with SILU, with $\alpha$=5%, 10% and 20% of the final steps using RELU. Note the loss spike when we switch the activation. The model rapidly recovers and converges to a regime where RELU is performing well while providing sparsity. This strategy needs to be combined with StochA to provide good models operating with RELU.
@@ 3.2 Stochastic activation: StochA
which amounts to randomly selecting between the RELU and SILU activations for $x<0$. If $x\geq 0$ we choose either $\Psi_{p}(x)=x$ or $\Psi_{p}(x)=\texttt{SILU}(x)$, see the baseline paragraph below.
@@ 3.3 Inference-time strategies and evaluation
RELU at test time.
@@
Stochastic activation at test time.
The following only applies to the StochA strategy: we evaluate the performance when if leverage the randomness at test time, i.e., in this case, we do not use RELU. This choice has two interests:
@@ 4.1 Experimental setting
We employ two types of benchmarks for zero or few-shot evaluation, which we describe in more detail in Appendix C and Table 6. The first type is code generation tasks: HumanEval+ (Liu et al., 2023) and MBPP (Chen et al., 2021). The second type consists of common sense and general reasoning: HellaSWAG(Zellers et al., 2019), ARC(Clark et al., 2018), PIQA (Bisk et al., 2020), OBQA (Mihaylov et al., 2018), WinoGrande (Sakaguchi et al., 2020), NQ (Kwiatkowski et al., 2019), RACE (Lai et al., 2017), TQA (Joshi et al., 2017) and GSM8K (Cobbe et al., 2021).
@@ 4.2 Performance analysis of Swi+FT and StochA with RELU at inference time
Table 1: Losses: train is computed over the last 500 steps of the training loss of LM1.5B, val is measured after training on a different set set of text and code using the RELU activation or StochA, i.e., the same activation used at train time (possibly deterministic). If Swi+FT is enabled, we switch to RELU for the last $5\%$ steps. : for the deterministic baselines SILU, S-R+ and S+R-, we do the inference with the same activation used at train-time (not RELU).
@@
Cross-entropy performance.
Table 1 provides the impact on the training and validation losses of multiple choices with the LM1.5B and LM3B models. We observe that the training loss using stochastic activation at train time is lower than that obtained with RELU. However the validation entropy is not competitive per se, due to the remaining train-inference discrepancy of activation. This is solved by Swi+FT: switching to the RELU activation function and fine-tuning for the last 5% or 10% steps percentage of the training steps drastically boosts the test-time inference. These results outperform the results obtained with regular RELU training, while using the same activation at test time.
@@
The activation sparsity can be exploited to avoid fetching 90% of the matrices $W_{1}$ and $W_{3}$ of the FFN (see Eq. 5 ands detailled sparsity rate in Appendix B). This has a direct benefit on CPU, see Figure 3: the 90% sparsity provides a 65% speedup. On GPU the additional challenge is to make the computation sufficiently predictable to balance the load between CUDA threads.
Figure 3: Total inference time for 1 token on CPU, as a function of the activation sparsity, with a LM3B model trained with Swi+FT. The “other operations" include the attention layers (they are not dominant because the generation is limited to 200 tokens), the normalization and the execution overheads. At 90% sparsity the speedup is $\times$1.65. The timings are measured on a single core of a Xeon 8462Y+ machine.
Complementary of StochA and Swi+FT with RELU finetuning
Figure 4 shows the training loss when using StochA and Swi+FT jointly. When switching the activation to RELU, we observe a spike in the loss, but the optimization rapidly recovers, and then converges to a model having a better performance than the one produced when training with RELU from scratch. In contrast, in the case where we employ Swi+FT alone (Figure 2), fine-tuning with RELU after pretraining with SILU is not enough to obtain performance improvements. This shows that both approaches are complementary.
@@
Figure 4: Training loss with Swi+FT and StochA: [S|R]-S+ activation with $p=0.5$ for $\alpha$$5\%$, $10\%$ and $20\%$, relative to RELU and SILU. Note that this plot this is zoomed in relative to Figure 2.
The impact of $\alpha$
can also be seen in Figure 4, where we compare the final loss reported for different fine-tuning values. In that case, setting $\alpha=0.1$ is a good trade-off since the corresponding loss is lower than the RELU loss.
@@ 4.3 Performance on downstream tasks
Table 2 reports the results for the standard code generation, common sense and general reasoning benchmarks detailed in Appendix C. We consider multiple StochA models using few-shot or zero shot prompting, see Table 6 in the appendix for more details for each case. Observe first that the model with SILU is significantly better that a regular model with RELU. However, our models trained with StochA are slightly better or on par with SILU: either the model fine-tuned with Swi+FT and using RELU at inference time, or even the model that uses StochA at test time.
@@
baselines (a) Swi+FT (b) StochA baselines (a) Swi+FT (b) StochA
@@
Table 2: Performance per benchmark of the RELU and SILU (a) baselines for LM1.5B and LM3B compared to (b) models with StochA and Swi+FT at train time and RELU at test time, and (c) models with StochA at train and test time. We use the model with the best perplexity on val namely $p=0.3,\alpha=0.05$ for Swi+FT and $p=0.5$ for StochA.
@@
Performance when varying $\boldsymbol{\alpha}$ with Swi+FT.
Figure 5 showcases that we can slightly surpass the SILU baseline if we first use a stochastic activation function during the LM1.5B model training and then switch to the RELU activation function for the last $\alpha\%$ of the training steps, for $\alpha\in\{5\%,10\%,20\%\}$. The best performance is typically obtained with $\alpha=5\%$ or $\alpha=10\%$ for the LM1.5B model.
@@ 4.4 Exploiting StochA at test time
Effectiveness of StochA a test time.
@@
Figure 6: Comparison of ways to generate diverse responses in Q&A benchmarks: by varying the softmax temperature or by varying the [S|R]-S+ pp. The multiple generations are scored by normalized log likelihoods and the best generation is evaluated with an F1 score (y-axis).
Diversity of generations ablations.
Figure 6 shows how stochastic generation compares to temperature sampling to generate diverse outputs, see Section D for examples. We plot these on TQA and NQ, whose results are least noisy. For the best settings, the curves are increasing, which means that (1) the generations are indeed diverse and (2) that the normalized log likelihoods are a suitable scoring function. The StochA activations yield a higher performance than vanilla temperature sampling on NQ but are sub-par for TQA.
@@ 5 Conclusion
This paper has introduced a novel stochastic activation that preserves the performance of a non-sparse activation, such as SILU, while better adjusting to the behavior of a sparse one, such as RELU, at test time. This improves the inference times for the FFN layers of a transformer, translating into a speedup of typically $\times 1.65$ for the FFN processing on CPUs while almost preserving the accuracy of the non-sparse SILU activation. Finally, we explore how stochastic activations can be leveraged at test time to improve diversity in model generations.
@@
In this appendix, we add some details that did not fit in the main paper. Appendix A gives some details about the models used in the paper. Appendix B gives additional details about the sparsity of the activations produced by the models that we use. Appendix C gives the references for all the benchmarks used in the experiments. The final Appendix D gives some examples of multiple answers generated by StochA.
@@ Appendix B Sparsity Analysis
baselines SILU SILU 00000.0002
RELU RELU 94.8
Swi+FT SILU RELU 0.05 79.9
@@
Table 5: We report the rate of the 0-valued activation after the SILU (not sparse) and RELU activation, when training with the LM1.5B model with the following activations: RELU, SILU, Swi+FT fine-tuning, and when using StochA with Swi+FT for varying values $p$.
Pre-activation inputs distribution dead neurons in the $W_{1}$ layer
@@
In Table 5, we report the sparsity rates resulting from the gating in the FFN layer, when using the following activations: RELU, SILU, Swi+FT fine-tuning, and when using StochA with Swi+FT for varying values $p$. In all cases, except for the SILU baseline, we use RELU at test time. We observe that Swi+FT fine-tuning with RELU brings back a lot of sparsity (80%), yet the model is not competitive. The combination of StochA with Swi+FT leads to models offering a high sparsity degree by using RELU at test time and simultaneously achieving competitive performance, see Tables 1 and 2.
@@
In Figure 7 (left), we observe that $P(X\leq 0)$ is the highest for RELU since most of its mass is concentrated around zero as in Table 5. For $x<0$, the probability $P(X\leq x)$ is lower for SILU than for both Swi+FT and when we combine StochA with Swi+FT for different values of $p$. This shows that combining StochA with Swi+FT $P(X\leq x)$ enables us to control how much mass we assign to negative values of $x$ depending on $p$.
The sudden jump in the CDF for RELU indicates that there are many activations that are exactly 0 on input of RELU. When investigating the rows of linear layer $W_{1}$ (Eq 5) in Figure 7 (right), it turns out a large fraction of these are near 0 (i.e. their norm is less than 1/1000 the average row norm of the matrix). This means that these rows are unused: they are “dead neurons”. Fortunately, with our Swi+FT approach, that uses SILU at pre-training, these dead neurons disappear. This explains why our approach can obtain significantly better results with RELU than the vanilla training of the model with the RELU activation employed from scratch.
Appendix C Benchmarks

### v2

@@ 1 Introduction
The majority of the LLM parameters are in the Feed-Forward Network (FFN) layers, where they primarily serve to store and recall information from the training data. FFNs are two linear layers separated by an activation function, and sometimes an additional linear layer that serves as a gating operation. The activation is a non-linear function from $\mathbb{R}$ to $\mathbb{R}$. In this context, the choice of activation function plays a crucial role for both the model’s expressivity and efficiency. The simplest choice of activation is RELU (Rectified Linear Unit), that allows positive values to pass through and forces negative inputs to zero. RELU is sparsity-inducing, since, on average, half of its outputs are zero (in practice significantly more). Within a two-layer Multilayer Perceptron, this means that inference on the second layer is a matrix-sparse vector multiplication, hence, it can be implemented with fewer FLOPs than a matrix-dense vector multiplication. Note that effectively improving the runtime with this sparsity pattern remains challenging.
@@
In this work, we consider two ways to approach this problem. The first approach is activation fine-tuning, denoted by Swi+FT: we pre-train the model with an activation that facilitates efficient large language model optimization, then, we change the activation to RELU and adapt the model by fine-tuning it further. Our second approach, referred to as StochA (stochastic activations), is a novel technique that randomly selects between multiple activations, either at train or test time. Both approaches allow models to benefit from the superior optimization properties of SILU. These hybrid strategies combine the best of both worlds – maintaining high model performance while unlocking the computational efficiency of sparse activations.
@@
  • We introduce two strategies that employ different activation functions at training and inference time, namely Swi+FT and StochA. Both are complementary and make it possible to use activations at inference time that differ from those employed during pre-training.
@@
  • We show that stochastic activations, when used at inference time, provide an alternative way to generate diverse sequences compared to traditional temperature sampling or other variants.
@@ 2 Related work
Activations are at the core of deep learning, they are fundamental to enable deep learning models to go beyond linear function with limited expressivity. While early neural network architectures were inspired by logistic regression, such as sigmoidal and tanh activations, many activation functions have been evaluated for the FFN layers of transformers. Vaswani et al. (2017) used RELU (Glorot et al., 2011) initially. However, using the RELU activation function leads to some neurons getting stuck in the negative region. As a consequence, models stop learning entirely, since the gradient is zero for negative inputs, and their weights do not get updated. In contrast, Touvron et al. (2023) used SILU as the activation function for the FFN layers of the transformer for the first Llama models. Shazeer (2020) discusses the benefits of SWIGLU, which consists of SILU with gating. There exist many other activation functions such as the Gaussian Error Linear Unit (GELU) (Hendrycks & Gimpel, 2016), Scaled Exponential Linear Unit (SELU) (Klambauer et al., 2017), Swish (Ramachandran et al., 2018) and gated SILU, among others.
@@
Lee et al. (2022) propose the Adaptive Swish(ASH) activation, which uses stochastic sampling of the top-k percentile elements. It generalizes the Swish (Ramachandran et al., 2018), which uses adaptive thresholding to select the values in the top percentiles and is set to zero otherwise. It is an example of a stochastic activation.
@@
LayerDrop (Fan et al., 2019) randomly drops entire layers during training, hence, it encourages the model to be robust to missing layers. At inference time, some layers can be pruned, trading off speed and accuracy as needed. While the method does not make the model sparse in the usual sense, it induces structured sparsity in the computation graph during training. Other works also introduce structured dropout variants such as DropBlock (Ghiasi et al., 2018), Bayesian dropout (Gal & Ghahramani, 2016), or Beit (Bao et al., 2021) and masked-autoencoder (He et al., 2022) in computer vision, among others.
@@ 3.1 Fine-tuning with RELU: Swi+FT
Figure 2: Swi+FT: Training loss. Most of the training is carried out with SILU, with $\alpha*100\%$=5%, 10% and 20% of the final steps using RELU. Note the loss spike when we switch the activation. The model rapidly recovers and converges to a regime where RELU is performing well while providing sparsity.
@@ 3.2 Stochastic activation: StochA
which corresponds to randomly selecting between the RELU and SILU activations for $x<0$. If $x\geq 0$ we choose $\Psi_{p}(x)=x$ or $\Psi_{p}(x)=\texttt{SILU}(x)$, see the baselines description below.
@@ 3.3 Inference-time strategies and evaluation
RELU at test time
@@
Stochastic activation at test time
The following only applies to the StochA strategy: we evaluate the performance by leveraging the randomness at test time, i.e., in this case, we do not use RELU. This choice is interesting for two reasons:
@@ 4.1 Experimental setting
We employ two types of benchmarks for zero or few-shot evaluation, which we describe in more detail in Appendix G and Table 8. The first type is code generation tasks: HumanEval+ (Liu et al., 2023) and MBPP (Chen et al., 2021). The second type consists of common sense and general reasoning: HellaSWAG(Zellers et al., 2019), ARC(Clark et al., 2018), PIQA (Bisk et al., 2020), OBQA (Mihaylov et al., 2018), WinoGrande (Sakaguchi et al., 2020), NQ (Kwiatkowski et al., 2019), RACE (Lai et al., 2017), TQA (Joshi et al., 2017) and GSM8K (Cobbe et al., 2021).
@@ 4.2 Performance analysis of Swi+FT and StochA with RELU at inference time
Table 1: The train loss is computed over the last 500 steps of the training of LM1.5B, the val loss is measured after training, on a different set of text and code, using the RELU activation or StochA, i.e. the same activation used at train time (possibly deterministic). If Swi+FT is enabled, we switch to RELU for the last $5\%$ steps. : for the deterministic baselines SILU, S-R+ and S+R-, we do the inference with the same activation used at train-time (not RELU).
@@
Cross-entropy performance
Table 1 summarizes the impact on the training and validation losses of multiple choices with the LM1.5B and LM3B models. We observe that the training loss using stochastic activation at train time is lower than that obtained with RELU. However the validation entropy is not competitive per se, due to the remaining train-inference discrepancy of activation. This is solved by Swi+FT: switching to the RELU activation function and fine-tuning for the last 5% or 10% steps of the training steps drastically boosts the test-time inference. These results outperform the results obtained with regular RELU training, while using the same activation at test time.
@@
[…]
CPU (one core of Xeon 8462Y+) GPU (H100 80GB HBM3)
Refer to caption Refer to caption
Figure 3: Inference times for 1 token, as a function of the activation sparsity, with a LM3B model trained with Swi+FT. For the CPU we indicate the total inference time, including “other operations": the attention layers (these are not dominant because the generation is limited to 200 tokens), the normalization and the execution overheads. At 90% sparsity the speedup is $\times$1.65. For the GPU, we measure only the FFN inference time. Note that the impact on inference is less important because at this speed and model scale the overheads are dominant. We also indicate the minimum runtime that could be obtained given the GPU’s memory bandwidth (roofline model).
Absence of dead neurons with Swi+FT
Figure 7(left) shows a sudden jump in the RELU CDF, which indicates there are many activations that are exactly zero on input of RELU. The RELU plot in Figure 7(right) shows that a large fraction of the rows of the linear layer $W_{1}$ (Eq 5) are near zero (i.e. their norm is less than 1/1000 the average row norm of the matrix). This means that these rows are unused: they are “dead neurons”. In contrast, the Swi+FT plot in Figure 7(right) we observe that using SILU at pre-training prevents these dead neurons appearing. This explains why our approach obtains significantly better results with RELU than the vanilla training of the model using the RELU from scratch.
Complementarity between StochA and Swi+FT when fine-tuned with RELU
Figure 4 shows the training loss when using StochA and Swi+FT jointly. When switching the activation to RELU, we observe a spike in the loss, but the optimization rapidly recovers, and converges to a model that has better performance than the one trained with RELU from scratch. In contrast, in the case where we employ Swi+FT alone (Figure 2), fine-tuning with RELU after pretraining with SILU is not enough to obtain performance improvements. As a consequence, the StochA and Swi+FT approaches are complementary.
@@
Figure 4: Training loss with Swi+FT and StochA: [S|R]-S+ activation with $p=0.5$ for $\alpha*100\%$$5\%$, $10\%$ and $20\%$, relative to RELU and SILU. This shows that the Swi+FT strategy needs to be combined with StochA to provide good models operating with RELU compared to finetuning SILU with RELU alone (Figure 2). This plot this is zoomed in relative to Figure 2.
Impact of fine-tuning with RELU for the last $\alpha*100\%$ steps
In Figure 4, we compare the final loss reported for different values of $\alpha$. In that case, setting $\alpha=0.1$ is a good trade-off since the corresponding loss is lower than the RELU loss.
Continuous pre-training with StochA and Swi+FT
Starting from a pre-trained LM1.5B model with SILU, we can leverage either StochA or Swi+FT methods for continuous pretraining (CPT), see Appendix E for experimental details. In such case, the best strategy is to use Swi+FT, i.e., simply fine-tune with RELU the model initially trained with SILU to obtain the best trade-off between sparsity and loss (Table 6). This pattern is the opposite of what we observe for LMs trained from scratch, where the StochA and StochA +Swi+FT offer the best trade-off between sparsity and loss (Table 1).
Generalizability of StochA and Swi+FT for any pair of activations
In Appendix F, we explain how our methods can be generalized for any pair of activation functions and include an illustrative example with the (TANH,RELU) activations.
@@ 4.3 Performance on downstream tasks
Table 2 reports the results for standard code generation, common sense and general reasoning benchmarks detailed in Appendix G. We consider multiple StochA models using few-shot or zero shot prompting, see Table 8 in Appendix G for more details. The model with SILU (topline) is significantly better that a regular model with RELU. However, our models trained with StochA are slightly better or on par with SILU: either the model fine-tuned with Swi+FT and using RELU at inference time, or even the model that uses StochA at test time.
@@
(a) baselines (b) Swi+FT (c) StochA (a) baselines (b) Swi+FT (c) StochA
@@
Table 2: Performance per benchmark of (a) RELU and SILU baselines for LM1.5B and LM3B compared to (b) models with StochA +Swi+FT at train time and RELU at test time, and (c) models with StochA at train and test time. We use the model with the best perplexity on val namely $p=0.3,\alpha=0.05$ for Swi+FT and $p=0.5$ for StochA. The average performance for the topline (SILU) is in bold, second best (StochA +Swi+FT) is underlined.
@@
Performance when varying $\boldsymbol{\alpha}$ with Swi+FT
Figure 5 shows that we can slightly surpass the SILU baseline average performance if we first use a stochastic activation function during the LM1.5B model training and then switch to the RELU activation for the last $\alpha*100\%$ of the training steps, for $\alpha\in\{0.05,0.1,0.2\}$. The best performance is obtained with $\alpha=0.05$ or $\alpha=0.1$ for the LM1.5B model.
@@ 4.4 Exploiting StochA at test time
Effectiveness of StochA a test time
@@
(b) Average type-token ratio:
Refer to caption Refer to caption
Figure 6: Comparison of diverse response generation methods in Q&A benchmarks. (a) F1 scores when varying the temperature with SILU or [S|R]-S+ $p$ with temp=0=0, the generations are scored by normalized log likelihoods. (b) TTR when varying the temperature in temperature sampling with SILU (solid lines) vs [S|R]-S+ for different values of $p$ (dashed lines).
Diversity of generated sequences ablation
Figure 6 shows that leveraging the StochA activations stochasticity (combined with standard temperature sampling) generates more diverse sequences compared to using SILU with standard temperature sampling. See Appendix C for the experimental setup and Appendix H for qualitative examples of such generations.
In Figure 6 (a), the curves are increasing, which means that (1) we obtain diverse generations and (2) that the normalized log-likelihood is a suitable scoring function. Specifically, for the NQ task, the StochA activation with $p=0.7$ and temperature zero yields consistently higher performance than standard temperature sampling with the SILU activation (for temperatures in $\{0.1,0.2,0.5\}$). The StochA activation with $p=0.3$ is also competitive for the number of generations up to five. In contrast, the results are subpar for the TQA task: standard temperature sampling with SILU activation yields higher performances for all number of generations for temperatures in $\{0.1,0.2\}$.
In Figure 6 (b), we observe that all the models that use StochA activation functions have consistently higher average TTR for both tasks, for each temperature in $\{0,0.1,0.2,0.3\}$ compared to standard temperature sampling with SILU. This finding further confirms that the text diversity of generations is much better for these models.
Token-level diversity ablation
In Appendix D, we highlight that using StochA at inference time is equivalent to applying Monte Carlo dropout (Gal & Ghahramani, 2016) on the negative side of the input after activation. We then leverage this token level diversity at inference time by doing multiple forward passes per generated token using greedy decoding. As observed in the sequence level ablation, more diverse tokens are generated. However, we did not observe that this resulted in better performance if the maximum probability or majority vote selection strategies were used.
@@ 5 Conclusion
This paper introduces a novel stochastic activation that preserves the performance of a non-sparse activation, such as SILU, while better adjusting to the behavior of a sparse one, such as RELU, at test time. This improves the inference times for the FFN layers of a transformer, translating into a speedup of typically $\times 1.65$ for the FFN processing on CPUs while almost preserving the accuracy of the non-sparse SILU activation. Finally, we explore how stochastic activations can be leveraged at test time to improve diversity in model generations.
@@
In this appendix, we add some extra details to complement the main paper. Appendix A gives architectural details for the models used in the paper. Appendix B gives additional details of the sparsity produced by the models that use the StochA and Swi+FT activations. Appendix C provides the experimental details for the diversity of generations ablation. Appendix D provides the experimental details for the multiple forward passes per generated token ablation. Appendix E gives some details for the CPT experiment with StochA and Swi+FT. Appendix F discusses how to generalize the StochA and Swi+FT methods for any pair of non-sparse and sparse activations. Appendix G gives the references for all the benchmarks used in the experiments. The final Appendix H gives some examples of multiple answers generated by StochA.
@@ Appendix B Sparsity Analysis
baselines SILU SILU - 00000.0002
RELU RELU - 94.8
Swi+FT SILU RELU - 0.05 79.9
@@
Table 5: We report the rate of the 0-valued activation after the SILU (not sparse) and RELU activation, when training with the LM1.5B model with the following activations: RELU, SILU, Swi+FT fine-tuning, and when using StochA with Swi+FT for varying values of $p$.
Pre-activation inputs distribution Presence of dead neurons in the $W_{1}$ layer
@@
In Table 5, we report the sparsity ratios resulting from the gating in the FFN layer, when using the following activations: RELU, SILU, Swi+FT fine-tuning, and when using StochA with Swi+FT for varying values $p$. In all cases, except for the SILU baseline, we use RELU at test time. We observe that Swi+FT fine-tuning with RELU brings back a lot of sparsity (80%), yet the model is not competitive. The combination of StochA with Swi+FT leads to models offering a high sparsity degree by using RELU at test time and simultaneously achieving competitive performance, see Tables 1 and 2.
@@
In Figure 7 (left), we observe that $P(X\leq 0)$ is the highest for RELU since most of its mass is concentrated around zero as in Table 5. For $x<0$, the probability $P(X\leq x)$ is lower for SILU than for both Swi+FT and when we combine StochA with Swi+FT for different values of $p$. This shows that combining StochA with Swi+FT enables us to control how much mass we assign to negative values of $x$ depending on $p$.
Appendix C Diversity of generations ablation experimental setup
We use the LM3B models trained with the StochA activation [S|R]-S+ with $p$ in $\{0.3,0.5,0.7,0.85\}$. At inference time, the probability $p$ is set to be the same as the one during training. We generate twenty answers for the NQ and TQA generation tasks, compute the Ff1 and Type-token ratio (TTR) metrics of $n$ generated answers (samples) and average over all datapoints for a given number of samples $n$ in $\{2,3,4,5,10,20\}$.
In Figure 6 (b) we report the type-token ratio metric (Johnson, 1944), a simple and widely used metric in linguistics and natural language processing to measure lexical diversity in a text. In this context, a type consists of a unique word in the text and a token consists of any word occurrence in the text (including repetitions). The Type-Token Ratio is then TTR=Number of Types / Number of Tokens.
Appendix D Multiple forward passes per generated token
If we use the StochA activation function, it is equivalent to doing Monte Carlo dropout (Gal & Ghahramani, 2016) on the negative side of the input after activation: it is set to zero with probability 1p1-p. For this reason, at test time, if we run multiple forward passes, this naturally creates a random output after activation which is either zero or SILU on the negative side of the inputs after activation.
In this setup, at generation time, after performing the multiple forward passes, we select the next token according to the following strategies:
Max probability: we select the token that has the highest probability.
Majority vote: we select the next token based on the token which appears the most often across the multiple passes.
Results. We use LM3B models trained with the [S|R]-S+ activations with $p$ in $\left\{0.3,0.5,0.7,0.85\right\}$. We vary the number of forward passes in $\left\{5,10,15,20,30\right\}$. We report the average performance of the generation tasks nq, tqa, mbpp and human eval plus and greedy decoding. Our findings indicate that even though we observe more diversity at the token level, the majority vote and maximum probability strategies do not perform better than the baseline, where we have a single forward pass as we increase the number of forward passes.
Appendix E Continuous Pretraining LM1.5B models
We can leverage the StochA and Swi+FT methods for continuous pretraining (CPT) language models, besides using these activations to train LMs from scratch. In this experiment, we start from the LM1.5B models pretrained with the SILU activation. During CPT, we use i) Swi+FT ii) StochA with $p=0.3$ or $p=0.5$ or iii) StochA and Swi+FT with $p=0.5$ and $\alpha=0.1$. We use the same number of steps used in the pretraining experiments: 45,629 steps. We compare the training losses against the baseline model that keeps SILU during CPT.
pretrain CPT inference StochA: p sparsity (%) train loss
baseline SILU SILU SILU - 0.0002 2.076
Swi+FT SILU RELU RELU - 90.96 2.086
StochA SILU StochA RELU 0.5 41.02 2.095
SILU StochA RELU 0.3 57.93 2.095
StochA +Swi+FT SILU StochA +Swi+FT RELU 0.5 81.61 2.096
Table 6: We report the rate of the 0-valued activations and training losses when doing CPT with the LM1.5B model with the following activations: RELU, SILU, Swi+FT,StochA and when using StochA with Swi+FT for varying values $p$.
Appendix F StochA and Swi+FT for (TANH, RELU) activations
Our StochA and Swi+FT methods can be generalized, starting with any pair of non-sparse, sparse activations, in the following way:
  1. 1.
    StochA: the non-sparse (activation which takes non-zero values on the negative side) and sparse activation pair can be used as a stochastic activation function that takes the form of either of the activations on the positive side, and, on the negative side, it either takes the non-zero value of the non-sparse variant with probability $p$ or is set to zero with probability 1p1-p.
  2. 2.
    Swi+FT: we start with an LM with a non-sparse activation and train it for $(1-\alpha)\%$ steps. Successively, we finetune the LM with the sparse activation for $\alpha\%$ of the final steps.
To illustrate how the StochA and Swi+FT methods are generalizable, we use the pair of (TANH, RELU) activations. We pretrain a LM1.5B with TANH as the non-sparse activation and RELU as the sparse one for 1) and 2). We compare these versus the losses to the corresponding baselines: using only the TANH or only the RELU activations, respectively. We denote the activation for 1) as Tanh+FT with parameter $\alpha$ and for 2), with [T|R]-T+ with parameter $p$.
We observe that there is barely any gap between the train loss of the non-sparse (TANH, bolded) and the sparse (RELU, in red) activations, hence, there is no trade-off between sparsity and performance, in contrast to the (SILU, RELU) case.
Training activation LM1.5B
$x<0$ $x>0$ $p$ train val val
Activation pp 1p1-p Swi+FT RELU StochA
TANH TANH TANH - 2.133 2.155
RELU RELU RELU - 2.140 2.162
Tanh+FT TANH TANH - 2.193 2.224
[T|R]-T+ TANH RELU TANH 0.3 2.368 10.803 2.421
[T|R]-T+ TANH RELU TANH 0.5 2.314 10.997 2.359
Table 7: Losses: train is computed over the last 500 steps of the training loss of LM1.5B, val is measured after training on a different set of text and code using the RELU activation or StochA, i.e., the same activation used at train time (possibly deterministic). If Swi+FT is enabled, we switch to RELU for the last $10\%$ steps.
Appendix G Benchmarks